# Probabilistic Analysis for Scalogram Ridges in Signal Processing

### v2

@@ Abstract
While ridges in the scalogram, determined by the squared modulus of analytic wavelet transform (AWT), is a widely accepted concept and utilized in nonstationary time series analysis, their behavior in noisy environments remains underexplored. Our object is to provide a theoretical foundation for scalogram ridges by defining ridges as a potentially set-valued random process connecting local maxima of the scalogram along the scale axis and analyzing their properties when the signal fulfills the adaptive harmonic model and is contaminated by stationary Gaussian noise. In addition to establishing several key properties of the AWT for random processes, we investigate the probabilistic characteristics of the resulting random ridge points in the scalogram. Specifically, we establish the uniqueness property of the ridge point at individual time instances and prove the upper hemicontinuity of the ridge random process. Furthermore, we derive bounds on the probability that the deviation between the ridges of noisy and clean signals exceeds a specified threshold, and these bounds depend on the signal-to-noise ratio. To achieve these ridge deviation results, we derive maximal inequalities for the complex modulus of nonstationary Gaussian processes, leveraging classical tools such as the Borell-TIS inequality and Dudley’s theorem, which might be of independent interest.
@@
adapted harmonic model; analytic wavelet transform; Borell-TIS inequality; circularly symmetric Gaussian process; maximum inequality; ridge; ridge deviation.
@@ 1 Introduction
Ridge analysis plays a vital role in connecting time-frequency analysis [daubechies1992ten, flandrin1998time] and signal processing [mallat1999wavelet], particularly for capturing dynamic information. Traditionally, ridges in time-frequency representations (TFRs) are interpreted ad hoc as sequences of local maxima along the frequency or scale axis, often visualized as continuous curves over mountain-shaped structures in the TFR. See Figure 1. This continuity is central to definitions for adapted harmonic signals [lilly2010analytic] and underlies many ridge extraction algorithms [carmona1999multiridge, carmona1997characterization, rankine2007if, Colominas_Meignen_Pham_2020, iatsenko2016extraction, Laurent_Meignen_2021, Legros_Fourer_2021, meignen2017demodulation, Ozkurt_Savaci_2005, su2024ridge, Zhu_Zhang_Gao_Li_2019]. Ridges encode instantaneous frequency (IF), making them foundational for various applications [flandrin1998time, mallat1999wavelet]. Despite widespread use and the development of many algorithms, the statistical properties of ridges and algorithmic accuracy under noise remain underexplored. While the term “ridge” also appears in other fields such as statistics and image analysis [hall1992ridge, eberly2012ridges], its meaning in TFRs is distinct. In what follows, we focus exclusively on ridges in TFRs.
@@
Following this framework, in Step I, we model the noisy signal as $Y=f+\Phi$, where $f$ fulfills the adaptive harmonic model (AHM) [2011synchrosqueezed] with possibly multiple oscillatory components having well-separated IFs, and $\Phi$ is modeled as a stationary Gaussian random process with potential long- or short-range dependence. We analyze the TFR of $Y$ obtained using the continuous wavelet transform with an analytic mother wavelet, referred to as the analytic wavelet transform (AWT), which is a complex-valued function $W_{Y}:(t,s)\to\mathbb{C}$, where $t\in\mathbb{R}$ is time and $s>0$ is scale, whose squared modulus $S_{Y}:(t,s)\to\mathbb{R}_{+}\cup\{0\}$, the scalogram, describes the spectral energy distribution over time. We call the set $\{t\in\mathbb{R},s>0\}$ the time-scale domain. It is important to note that when the analysis centers on IF, researchers typically employ the AHM to facilitate the estimation of IFs by extracting energy-concentrated curves from the scalograms (or spectrograms) of noisy signals [carmona1997characterization, carmona1999multiridge, iatsenko2016extraction, stiles2004wavelet, su2024ridge].
In Step II, we examine the behavior of $W_{Y}$. When the IFs of $f$ are well-separated, the scalogram of $f$, denoted by $S_{f}$, encodes IF information through its local maxima along the scale axis [delprat1992asymptotic]. The AWT of the Gaussian noise process $\Phi$, denoted by $W_{\Phi}$, interacts with the AWT of $f$, denoted by $W_{f}$, via the relation $S_{Y}=|W_{f}+W_{\Phi}|^{2}$. How the random field $W_{\Phi}$ perturbs the local maxima in $S_{Y}$ is the main focus of this paper. By viewing $S_{Y}$ as an inhomogeneous random field, the study of its local maxima is naturally related to excursion set analysis [adler2010geometry], though the current theory does not fully account for the nonstationarity or the temporal coherence critical to ridge structure. Additional questions, such as how the local maxima of $S_{f}$, and thus of $S_{Y}$, are affected by spectral inference when the IFs of $f$ are close [delprat2002global] or what local maxima mean for the general $f\in L^{2}$ are important but deferred to future work.
@@
commonly known as the wavelet ridge [mallat1999wavelet] or wavelet skeleton [antoine2008two, pp. 14–18]. This definition extends naturally to the observed random process by
@@
with $s_{Y}$ referred to as the ridge of $Y$. Although widely used, the properties of $s_{Y}$ remain largely unexplored. A central challenge lies in the fact that as a nonlinear functional of the random field $W_{Y}$, $s_{Y}$ is a random process that may be discontinuous or multi-valued. When $f$ consists of multiple oscillatory components, each corresponds to a local maximum in $S_{f}$ [delprat1992asymptotic], and under sufficient IF separation, individual ridges can be defined by segmenting the time-scale domain. We explain this in detail later. We shall mention that if the IFs are not well separated or even crossover, ridge definitions become ambiguous, even without noise [delprat2002global, meignen2021study], and this is beyond the scope of this paper.
@@
Our contributions in this paper are multifaceted. First, we analyze the AWT, focusing on its regularity and a novel covariance structure of the scalogram of $\Phi$. These aspects are critical but underexplored aspects of statistical inference. Second, we propose a ridge definition for noisy signals suited to practical needs. Under mild conditions, we show that the set-valued random process $t\mapsto s_{Y}(t)$ is almost surely single-valued on any countable and dense subset of $\mathbb{R}$, and upper hemicontinuous on $\mathbb{R}$. This means that for any time $t$, as $t^{\prime}$ approaches $t$, the maxima of $S_{Y}(t^{\prime},\cdot)$ converge arbitrarily close to some of the maxima of $S_{Y}(t,\cdot)$. Based on these properties, we further discuss the local $C^{1}$-smoothness of $s_{Y}(t)$. Third, we analyze how $s_{Y}(t)$ deviates from $s_{f}(t)$. Since the scalogram involves a nonlinear modulus operator, analyzing $|s_{Y}(t)-s_{f}(t)|$ is difficult, particularly from the perspectives of time-frequency analysis and probability. This problem has received limited attention in the literature, with the exception of [sejdic2008quantitative], where the authors focused on a frequency-modulated signal contaminated with additive Gaussian white noise and provided approximations for $\mathbb{E}[s_{Y}(t)-s_{f}(t)]$ and $\textup{Var}(s_{Y}(t)-s_{f}(t))$. Our work extends this analysis to temporally correlated Gaussian noise models. In Theorems 4 and 5, we provide upper bounds on the probability that $|s_{Y}(t)-s_{f}(t)|$ exceeds a given threshold, where the bounds decay exponentially as the noise strength decreases, offering theoretical support for the numerical observations in Figure 6. These bounds are derived from new maximal inequalities for $|W_{\Phi}|$ and its derivative (Lemma 4). These extend classical maximal inequalities, such as the Borell-TIS inequality and Dudley’s theorem, to complex-valued Gaussian processes arising from the AWT of $\Phi$ and may be of independent interest because the classical Gaussian concentration inequalities [kuhn2023maximal], Dudley’s theorem, and the Borell-TIS inequality [talagrand2014upper, vershynin2018high] are formulated for real-valued Gaussian elements. Furthermore, Theorem 5 reveals how the spectral energy distribution of the noise influences the probability of ridge deviations, yielding insights that go beyond what can be obtained through the approximations of $\mathbb{E}[s_{Y}(t)-s_{f}(t)]$ and $\textup{Var}(s_{Y}(t)-s_{f}(t))$.
@@
The rest of the paper is organized as follows. In Section 2, we summarize the necessary material for AWT and present the definitions of the adapted harmonic model and the noise we consider. Section 3 is dedicated to AWT for Gaussian random processes, where we present both old and new results. In Section 4, we introduce a definition of the ridge that takes into account practical needs when the data is noisy. In Section 5, we state our main results about the probability of the occurrence of ridge deviation. Section 6 concludes the paper with a discussion of the results and future directions. The proofs of our main results and some technical lemmas are provided in Section 7. Finally, a list of frequently used symbols is summarized in Table 1 in the appendix.
@@
Figure 1: The top section illustrates a signal $f$, composed of two frequency-modulated sinusoids, along with its 3-dimensional and image representations in the time-scale domain. Here, the mother wavelet used has center frequency $\omega_{\psi}=80$ Hz, as defined in (16). Hence, the signal $f$’s IF lies approximately within the interval $\frac{\omega_{\psi}}{58},\frac{\omega_{\psi}}{10}]$ Hz. The bottom section presents the same for its noise-affected counterpart $Y$, with a signal-to-noise ratio of -8.98 dB. Here, $Y$ is the sum of $f$ and a sample path of the Gaussian process $\Phi$. For the single-component case, see Figure 8 in the appendix.
@@ 2 Preliminary
where $A_{m}(t)\in C^{1}(\mathbb{R})$ are positive functions representing the amplitude modulation and $\phi_{m}(t)\in C^{2}(\mathbb{R})$ are strictly monotonically increasing functions representing the phase so that $\phi^{\prime}_{m}(t)>0$ represents the IF. We impose two conditions on the AHM model. The first one is the slowly varying condition: $|A^{\prime}_{m}(t)|\leq\epsilon\phi^{\prime}_{m}(t)$ and $|\phi^{\prime\prime}_{m}(t)|\leq\epsilon\phi^{\prime}_{m}(t)$, where $\epsilon\geq 0$ is a small constant. The second one is the spectral separation condition when $M>1$:
@@
for all $t\in\mathbb{R}$ so that the IF separation condition $\phi^{\prime}_{m+1}(t)-\phi^{\prime}_{m}(t)\geq\delta(\phi^{\prime}_{m+1}(t)+\phi^{\prime}_{m}(t))$ for some $\delta\in(0,1)$ is satisfies. When $M=1$, we denote $A(t)=A_{1}(t)$ and $\phi(t)=\phi_{1}(t)$. We call $A_{m}(t)\cos(2\pi\phi_{m}(t))$ the $m$-th intrinsic mode type (IMT) function. While it is possible to consider more complicated models, such as those involving crossover IF [chen2023disentangling], IFs with insufficient separation [meignen2021study], or non-sinusoidal oscillation [su2024ridge], in this paper we focus on the AHM model defined above.
@@
for any $\Delta_{1},\Delta_{2}\in\mathcal{B}(\mathbb{R})$, where Leb is the Lebesgue measure on $\mathbb{R}$ and $\mathcal{B}(\mathbb{R})$ is the Borel $\sigma$-algebra on $\mathbb{R}$ [major1981lecture].
The analytic wavelet transform (AWT) of an $L^{2}$ function $f$, denoted by $\{W_{f}(t,s)\mid t\in\mathbb{R},s>0\}$, is defined as follows [daubechies1992ten]:
@@
Similar to (6) and (7), we denote the AWTs of $\Phi$ and $Y$ by $W_{\Phi}$ and $W_{Y}$, respectively, and the scalograms of $\Phi$ and $Y$ by $S_{\Phi}$ and $S_{Y}$, respectively. By the linearity of AWT, we know $W_{Y}=W_{f}+W_{\Phi}$. By (4) and the stochastic Fubini Theorem [pipiras2010regularization],
@@ 3 AWT for Gaussian random processes–some old and new results
When the noise is modeled as a stationary Gaussian process, the scalogram $S_{\Phi}$ at a given time is a nonstationary random process along the scale axis. Theorem 1(c) shows that for any $t\in\mathbb{R}$, the nonstationary process $S_{\Phi}(t,\cdot)$ is non-negatively correlated. Because this property is not found in existing literature, we provide its proof in Section 7.1.
@@
$\textup{($\mathrm{D}^{0}_{1}$-$\mathrm{D}^{0}_{2}$):}\ \underset{\lambda>1}{\sup}|\lambda^{p}\widehat{\psi}(\lambda)|<\infty,\ p=1,2;\ \ \$$\textup{($\mathrm{D}^{1}_{0}$-$\mathrm{D}^{1}_{2}$):}\ \underset{\lambda>0}{\sup}|\lambda^{p}D\widehat{\psi}(\lambda)|<\infty,\ p=0,1,2;$
@@
The set of wavelets that satisfy Assumption 2 includes some of the generalized Morse wavelets discussed in [lilly2010analytic, (14)] and the analytic wavelets derived in [holighaus2019characterization, (6)]. For instance, the generalized Morse wavelets, whose Fourier transform takes the form $\widehat{\psi}(\lambda)=a_{\beta_{1},\beta_{2}}\lambda^{\beta_{1}}e^{-\lambda^{\beta_{2}}}1_{[0,\infty)}(\lambda),\ \beta_{1}\geq 1,\ \beta_{2}>0$, where $a_{\beta_{1},\beta_{2}}$ is a normalizing constant. Another example is the Klauder wavelet, whose Fourier transform is given by $\widehat{\psi}(\lambda)=\lambda^{\alpha}e^{-\gamma\lambda}e^{i\beta\log\lambda}1_{[0,\infty)}(\lambda)$, where $\alpha\geq 1$, $\beta\in\mathbb{R}$, and $\gamma\in\mathbb{C}$ with $\textup{Re}(\gamma)>0$.
@@
Let us consider the stationary random process $\Phi$ with the generalized Linnik covariance function [lim2010analytic], defined as
@@
According to [lim2010analytic, Propositions 3.1 and 3.4], $p(\lambda)=\mathcal{O}(|\lambda|^{-(1+\gamma)})$ when $|\lambda|\rightarrow\infty$. For the case $H>1$, because $C_{\Phi}\in L^{1}(\mathbb{R})$, the spectral density function $p$ is continuous everywhere. For the case $0<H<1$, by [lim2010analytic, Corollaries 3.3 and 3.10],

### v3

# Probabilistic Analysis of Scalogram Ridges in Signal Processing

@@ Abstract
While ridges in the scalogram, determined by the squared modulus of analytic wavelet transform (AWT), is a widely accepted concept and utilized in nonstationary time series analysis, their behavior in noisy environments remains underexplored. Our objective is to provide a theoretical foundation for scalogram ridges by defining ridges as a potentially set-valued random process connecting local maxima of the scalogram along the scale axis and analyzing their properties when the signal fulfills the adaptive harmonic model and is contaminated by stationary Gaussian noise. In addition to establishing several key properties of the AWT for random processes, we investigate the probabilistic characteristics of the resulting random ridge points in the scalogram. Specifically, we establish the uniqueness property of the ridge point at individual time instances and prove the upper hemicontinuity of the ridge random process. Furthermore, we derive bounds on the probability that the deviation between the ridges of noisy and clean signals exceeds a specified threshold, and these bounds depend on the signal-to-noise ratio. To achieve these ridge deviation results, we derive maximal inequalities for the complex modulus of nonstationary Gaussian processes, leveraging classical tools such as the Borell-TIS inequality and Dudley’s theorem, which might be of independent interest.
@@
adaptive harmonic model; analytic wavelet transform; Borell-TIS inequality; circularly symmetric Gaussian process; maximum inequality; ridge; ridge deviation.
@@ 1 Introduction
Ridge analysis plays a vital role in connecting time-frequency analysis [11, 18] and signal processing [30], particularly for capturing dynamic information. Traditionally, ridges in time-frequency representations (TFRs) are interpreted ad hoc as sequences of local maxima along the frequency or scale axis, often visualized as continuous curves over mountain-shaped structures in the TFR. See Figure 1. This continuity is central to definitions for adaptive harmonic signals [26] and underlies many ridge extraction algorithms [5, 6, 39, 9, 22, 24, 25, 33, 35, 44, 48]. Ridges encode instantaneous frequency (IF), making them foundational for various applications [18, 30]. Despite widespread use and the development of many algorithms, the statistical properties of ridges and algorithmic accuracy under noise remain underexplored. While the term “ridge” also appears in other fields such as statistics and image analysis [19, 16], its meaning in TFRs is distinct. In what follows, we focus exclusively on ridges in TFRs.
@@
Following this framework, in Step I, we model the noisy signal as $Y=f+\Phi$, where $f$ fulfills the adaptive harmonic model (AHM) [10] with possibly multiple oscillatory components having well-separated IFs, and $\Phi$ is modeled as a stationary Gaussian random process with potential long- or short-range dependence. We analyze the TFR of $Y$ obtained using the continuous wavelet transform with an analytic mother wavelet, referred to as the analytic wavelet transform (AWT), which is a complex-valued function $W_{Y}:(t,s)\to\mathbb{C}$, where $t\in\mathbb{R}$ is time and $s>0$ is scale, whose squared modulus $S_{Y}:(t,s)\to\mathbb{R}_{+}\cup\{0\}$, the scalogram, describes the spectral energy distribution over time. We call the set $\{t\in\mathbb{R},s>0\}$ the time-scale domain. It is important to note that when the analysis centers on IF, researchers typically employ the AHM to facilitate the estimation of IFs by extracting energy-concentrated curves from the scalograms (or spectrograms) of noisy signals [6, 5, 22, 43, 44].
In Step II, we examine the behavior of $W_{Y}$. When the IFs of $f$ are well-separated, the scalogram of $f$, denoted by $S_{f}$, encodes IF information through its local maxima along the scale axis [12]. The AWT of the Gaussian noise process $\Phi$, denoted by $W_{\Phi}$, interacts with the AWT of $f$, denoted by $W_{f}$, via the relation $S_{Y}=|W_{f}+W_{\Phi}|^{2}$. How the random field $W_{\Phi}$ perturbs the local maxima in $S_{Y}$ is the main focus of this paper. By viewing $S_{Y}$ as an inhomogeneous random field, the study of its local maxima is naturally related to excursion set analysis [1], though the current theory does not fully account for the nonstationarity or the temporal coherence critical to ridge structure. Additional questions, such as how the local maxima of $S_{f}$, and thus of $S_{Y}$, are affected by spectral inference when the IFs of $f$ are close [13] or what local maxima mean for the general $f\in L^{2}$ are important but deferred to future work.
@@
commonly known as the wavelet ridge [30] or wavelet skeleton [3, pp. 14–18]. This definition extends naturally to the observed random process by
@@
with $s_{Y}$ referred to as the ridge of $Y$. Although widely used, the properties of $s_{Y}$ remain largely unexplored. A central challenge lies in the fact that as a nonlinear functional of the random field $W_{Y}$, $s_{Y}$ is a random process that may be discontinuous or multi-valued. When $f$ consists of multiple oscillatory components, each corresponds to a local maximum in $S_{f}$ [12], and under sufficient IF separation, individual ridges can be defined by segmenting the time-scale domain. We explain this in detail later. We shall mention that if the IFs are not well separated or even crossover, ridge definitions become ambiguous, even without noise [13, 32], and this is beyond the scope of this paper.
@@
Our contributions in this paper are multifaceted. First, we analyze the AWT, focusing on its regularity and a novel covariance structure of the scalogram of $\Phi$. These aspects are critical but underexplored aspects of statistical inference. Second, we propose a ridge definition for noisy signals suited to practical needs. Under mild conditions, we show that the set-valued random process $t\mapsto s_{Y}(t)$ is almost surely single-valued on any countable and dense subset of $\mathbb{R}$, and upper hemicontinuous on $\mathbb{R}$. This means that for any time $t$, as $t^{\prime}$ approaches $t$, the maxima of $S_{Y}(t^{\prime},\cdot)$ converge arbitrarily close to some of the maxima of $S_{Y}(t,\cdot)$. Based on these properties, we further discuss the local $C^{1}$-smoothness of $s_{Y}(t)$. Third, we analyze how $s_{Y}(t)$ deviates from $s_{f}(t)$. Since the scalogram involves a nonlinear modulus operator, analyzing $|s_{Y}(t)-s_{f}(t)|$ is difficult, particularly from the perspectives of time-frequency analysis and probability. This problem has received limited attention in the literature, with the exception of [42], where the authors focused on a frequency-modulated signal contaminated with additive Gaussian white noise and provided approximations for $\mathbb{E}[s_{Y}(t)-s_{f}(t)]$ and $\textup{Var}(s_{Y}(t)-s_{f}(t))$. Our work extends this analysis to temporally correlated Gaussian noise models. In Theorems 3 and 4, we provide upper bounds on the probability that $|s_{Y}(t)-s_{f}(t)|$ exceeds a given threshold, where the bounds decay exponentially as the noise strength decreases, offering theoretical support for the numerical observations in Figure 6. These bounds are derived from new maximal inequalities for $|W_{\Phi}|$ and its derivative (Lemma 4). These extend classical maximal inequalities, such as the Borell-TIS inequality and Dudley’s theorem, to complex-valued Gaussian processes arising from the AWT of $\Phi$ and may be of independent interest because the classical Gaussian concentration inequalities [23], Dudley’s theorem, and the Borell-TIS inequality [45, 46] are formulated for real-valued Gaussian elements. Furthermore, Theorem 4 reveals how the spectral energy distribution of the noise influences the probability of ridge deviations, yielding insights that go beyond what can be obtained through the approximations of $\mathbb{E}[s_{Y}(t)-s_{f}(t)]$ and $\textup{Var}(s_{Y}(t)-s_{f}(t))$.
@@
The rest of the paper is organized as follows. In Section 2, we summarize the necessary material for AWT and present the definitions of the adaptive harmonic model and the noise we consider. Section 3 is dedicated to AWT for Gaussian random processes, where we present both old and new results. In Section 4, we introduce a definition of the ridge that takes into account practical needs when the data is noisy. In Section 5, we state our main results about the probability of the occurrence of ridge deviation. Section 6 concludes the paper with a discussion of the results and future directions. The proofs of our main results and some technical lemmas are provided in Section 7. Finally, a list of frequently used symbols is summarized in Table 1 in the appendix.
@@
Figure 1: The top panel illustrates a signal $f$, composed of two frequency-modulated sinusoids, along with its 3-dimensional and image representations in the time-scale domain. Here, the mother wavelet used has center frequency $\omega_{\psi}=80$ Hz, as defined in (16). Hence, the signal $f$’s IF lies approximately within the interval $\frac{\omega_{\psi}}{58},\frac{\omega_{\psi}}{10}]$ Hz. The bottom panel presents the same for its noise-affected counterpart $Y$, with a signal-to-noise ratio of -8.98 dB. Here, $Y$ is the sum of $f$ and a sample path of the Gaussian process $\Phi$. The y-axis represents the scale variable $s$, which is dimensionless. The tick values $10,14,\ldots,58$ are scale values rather than physical frequencies. For a mother wavelet with central frequency $\omega_{\psi}=80$ Hz, a scale value $s$ corresponds to the frequency $\omega_{\psi}/s$ Hz. Thus, the tick values $10,14,\ldots,58$ correspond to the frequencies $80/10,80/14,\ldots,80/58$ Hz, respectively. For the single-component case, see Figure 8 in the appendix.
@@ 2 Preliminary
where $A_{m}(t)\in C^{1}(\mathbb{R})$ are positive functions representing the amplitude modulation (AM) and $\phi_{m}(t)\in C^{2}(\mathbb{R})$ are strictly monotonically increasing functions representing the phase so that $\phi^{\prime}_{m}(t)>0$ represents the IF. We impose two conditions on the AHM model. The first one is the slowly varying condition: $|A^{\prime}_{m}(t)|\leq\epsilon\phi^{\prime}_{m}(t)$ and $|\phi^{\prime\prime}_{m}(t)|\leq\epsilon\phi^{\prime}_{m}(t)$, where $\epsilon\geq 0$ is a small constant. The second one is the spectral separation condition when $M>1$:
@@
for all $t\in\mathbb{R}$ so that the IF separation condition $\phi^{\prime}_{m+1}(t)-\phi^{\prime}_{m}(t)\geq\delta(\phi^{\prime}_{m+1}(t)+\phi^{\prime}_{m}(t))$ for some $\delta\in(0,1)$ is satisfies. When $M=1$, we denote $A(t)=A_{1}(t)$ and $\phi(t)=\phi_{1}(t)$. We call $A_{m}(t)\cos(2\pi\phi_{m}(t))$ the $m$-th intrinsic mode type (IMT) function. While it is possible to consider more complicated models, such as those involving crossover IF [7], IFs with insufficient separation [32], or non-sinusoidal oscillation [44], in this paper we focus on the AHM model defined above.
@@
for any $\Delta_{1},\Delta_{2}\in\mathcal{B}(\mathbb{R})$, where Leb is the Lebesgue measure on $\mathbb{R}$ and $\mathcal{B}(\mathbb{R})$ is the Borel $\sigma$-algebra on $\mathbb{R}$ [29].
The analytic wavelet transform (AWT) of an $L^{2}$ function $f$, denoted by $\{W_{f}(t,s)\mid t\in\mathbb{R},s>0\}$, is defined as follows [11]:
@@
Similar to (6) and (7), we denote the AWTs of $\Phi$ and $Y$ by $W_{\Phi}$ and $W_{Y}$, respectively, and the scalograms of $\Phi$ and $Y$ by $S_{\Phi}$ and $S_{Y}$, respectively. By the linearity of AWT, we know $W_{Y}=W_{f}+W_{\Phi}$. By (4) and the stochastic Fubini Theorem [36],
@@ 3 AWT for Gaussian random processes–some old and new results
When the noise is modeled as a stationary Gaussian process, the scalogram $S_{\Phi}$ at a given time is a nonstationary random process along the scale axis. Proposition 1(c) shows that for any $t\in\mathbb{R}$, the nonstationary process $S_{\Phi}(t,\cdot)$ is non-negatively correlated. Because this property is not found in existing literature, we provide its proof in Section 7.1.
@@
$\textup{($\mathrm{D}^{0}_{1}$-$\mathrm{D}^{0}_{2}$):}\ \underset{\lambda>0}{\sup}|\lambda^{p}\widehat{\psi}(\lambda)|<\infty,\ p=1,2;\ \ \$$\textup{($\mathrm{D}^{1}_{0}$-$\mathrm{D}^{1}_{2}$):}\ \underset{\lambda>0}{\sup}|\lambda^{p}D\widehat{\psi}(\lambda)|<\infty,\ p=0,1,2;$
@@
Assumption 2 imposes mild smoothness and polynomial decay conditions on the Fourier transform of the analytic mother wavelet. Since the polynomial decay of the Fourier transform suppresses high-frequency components in the noise, these conditions ensure that the scalogram of the analytic wavelet transform of the stationary Gaussian noise process $\Phi$ admits sample paths that are twice continuously differentiable on the time-scale domain; see Proposition 2. The regularity is essential for establishing the local uniqueness and hemicontinuity properties of ridge curves (Theorems 1 and 2). In particular, the uniqueness of the ridge point for the noise-affected signal guarantees that the deviation of the ridge point from its noise-free counterpart is well defined. On the other hand, under the conditions $(\mathrm{D}^{0}_{1})$, $(\mathrm{D}^{1}_{0})$, and $(\mathrm{D}^{1}_{1})$, the continuity of the AWT of the Gaussian process $\Phi$ can be extended to the boundary of the time-scale domain by defining $W_{\Phi}(t,0)=0$. This boundary continuity is a key ingredient in our extension of Dudley’s theorem to $\mathbb{C}$-valued nonstationary Gaussian random processes indexed by scale and generated by the AWT (see the proof of Lemma 3).
@@
The set of wavelets that satisfy Assumption 2 includes some of the generalized Morse wavelets discussed in [26, (14)] and the analytic wavelets derived in [20, (6)]. For instance, the generalized Morse wavelets, whose Fourier transform takes the form $\widehat{\psi}(\lambda)=a_{\beta_{1},\beta_{2}}\lambda^{\beta_{1}}e^{-\lambda^{\beta_{2}}}1_{[0,\infty)}(\lambda),\ \beta_{1}\geq 1,\ \beta_{2}>0$, where $a_{\beta_{1},\beta_{2}}$ is a normalizing constant. Another example is the Klauder wavelet, whose Fourier transform is given by $\widehat{\psi}(\lambda)=\lambda^{\alpha}e^{-\gamma\lambda}e^{i\beta\log\lambda}1_{[0,\infty)}(\lambda)$, where $\alpha\geq 1$, $\beta\in\mathbb{R}$, and $\gamma\in\mathbb{C}$ with $\textup{Re}(\gamma)>0$.
@@
Let us consider the stationary random process $\Phi$ with the generalized Linnik covariance function [27], defined as
@@
According to [27, Propositions 3.1 and 3.4], $p(\lambda)=\mathcal{O}(|\lambda|^{-(1+\gamma)})$ when $|\lambda|\rightarrow\infty$. For the case $H>1$, because $C_{\Phi}\in L^{1}(\mathbb{R})$, the spectral density function $p$ is continuous everywhere. For the case $0<H<1$, by [27, Corollaries 3.3 and 3.10],